Determine (f o g)(x) and (g o f )(x) for the pair of functions given. Also specify the domain of (f o g)(x) and (g 0 f )(x).
f(x) = 5x and g(x) = 9x-3

Answer by robertb(5830)

(f o g)(x) = f(g(x)) = f (9x - 3) = 5(9x-3) = 45x - 15. Domain is the set of all real numbers.
(g o f)(x) = g(f(x)) = g(5x) = 9*5x - 3 = 45x - 3. Domain is the set of all real numbers.